OpenStudy (anonymous):

a country's population in 1994 was 140 million. In 1998 it was 146 million. Estimate the population in 2011 using the exponential growth formula. Round your answer to the nearest million. P=Ae^kt

OpenStudy (anonymous):

0.7!!!!!!!!!!!! THERE!

OpenStudy (anonymous):

what does 0.7 represent?

OpenStudy (wolf1728):

42 cookiethere isn't here anymore I'll get an answer for you.

OpenStudy (anonymous):

does .07 represent k?

OpenStudy (wolf1728):

I do not know what k represents Maybe 42cookie just typed anything I'll work on the answer

OpenStudy (campbell_st):

1st task is to find k so \[146 = 140 \times e^{4k}\] solve for k

OpenStudy (anonymous):

2.269?

OpenStudy (campbell_st):

no... divide both sides by 140\[\frac{146}{140} = e^{4k}\] then take the ln of both sides and divide the answer by 4 that will give you a value for the growth constant k... then use that with P = 140 and t = 17 to find the population in 2011

OpenStudy (anonymous):

I'm gettinf 0.167

OpenStudy (anonymous):

getting*

OpenStudy (wolf1728):

ln means natural log ln (146/140) = ln (e^4*k) 0.04196419914 = 4*k

OpenStudy (anonymous):

Couldn't figure it out but thanks!

OpenStudy (wolf1728):

k = 0.0104910498

OpenStudy (wolf1728):

Population 1998 = 140 mil * e^(0.0104910498 * 4) Population 1998 = 140 * 1.0428571429 Population 1998 = 146 million

OpenStudy (wolf1728):

Population 2011 = 140 mil * e^(0.0104910498 * 17) Population 2011 = 140 mil * e^(0.1783478466) Population 2011 = 140 mil * 1.1952410094 Population 2011 = 167,333,741 Yes, it's just that easy!!! LOL